# A Novel Intensity-Based Criterion for Determining Exceptional Point in PT-Symmetry Coupled Array of Optical Waveguides

### v1

@@ Abstract
In this study, we investigated the propagation pattern and the site-to-site correlation function in a PT-symmetric waveguide array with different input quantum states. Recognizing the stark difference in propagation pattern before and after the PT symmetry-breaking point, we have developed a novel, straightforward intensity-based criterion to determine the exceptional point (EP). This new criterion shows excellent agreement with those obtained by directly computing the Hamiltonian’s eigenvalues. Given the computational complexity of Hamiltonian diagonalization, our proposed criterion provides a highly efficient and valuable alternative for identifying the PT symmetry-breaking point.
@@ 1 Introduction
Over the past decade, discrete optical systems-particularly waveguide arrays-have played a crucial role in controlling the behavior of light [1]. Optical waveguide arrays consist of several single-mode waveguides arranged at specific distances. In this system, light is transferred between adjacent waveguides through evanescent coupling. The most effective method for manufacturing these waveguides is direct writing with femtosecond lasers [2].
@@
The ability to control various parameters within waveguides offers a significant advantage for investigating PT-symmetric systems. PT-symmetry in quantum mechanics was first explored in 1998, following the discovery of Hamiltonians that remain invariant under both parity and time-reversal operations [3]. Notably, these non-Hermitian Hamiltonians can exhibit entirely real spectra [4]. PT-symmetry was introduced into optics due to the similarity between the paraxial wave equation and the Schrödinger equation, with the refractive index serving as the optical equivalent of the potential [5]. A PT-symmetric system in optics can be realized by adjusting the real and imaginary components of the refractive index [6]. For PT-symmetry to hold in optics, the real part of the refractive index needs to be spatially even, while its imaginary part must be odd function of position [7]. Since the imaginary part of the refractive index defines gain and loss, a PT-symmetric system can be achieved in optics by tuning these parameters accordingly. In a two-component optical system, real eigenvalues arise when the gain in one component compensates for the loss in the other. In contrast, in the weak coupling regime, if the gain from one component cannot fully compensate for the loss in the other, the system exhibits complex eigenvalues. This phase transition from real to complex spectra occurs at the EP [8].
@@
A non-Hermitian Hamiltonian typically leads to complex eigenvalues, where the negative imaginary part represents a damping constant associated with the finite spectral linewidth of the resonance energy [9]. The real part of the eigenvalue determines this linewidth. In classical optical systems, the resonance frequencies and modes correspond to the eigenvalues and eigenstates of the Hamiltonian. Non-Hermiticity can arise due to absorption, gain, and radiation. When the Hamiltonian is non-Hermitian, i.e $\hat{H}\neq\hat{H}^{\dagger}$, its eigenstates lose their orthogonality. The non-orthogonality is maximized at the EP, where the eigenstates of the Hamiltonian coalesce, and they can no longer form a complete basis. In this scenario, the states become parallel to one another. These points are considered singularities within non-Hermitian systems. Near an EP, one or more eigenvalues and their corresponding eigenstates coalesce, leading to a phase transition. This transition, known as a PT-symmetry phase transition, results in the emergence of complex eigenvalues [8, 10].
@@
Thompson et al. [11] examined intensity and site-to-site correlations in disordered optical waveguide arrays under Anderson localization, using non-classical (squeezed) and classical-like (coherent) light. However, their system did not exhibit PT symmetry. In 2018, the PT-symmetry breaking point of two curved, closely spaced PT-symmetric waveguides, characterized by a small period, was numerically computed using COMSOL Multiphysics software [12]. The investigation of boundary effects on propagation dynamics and site-to-site correlations within optical waveguide arrays in the quantum regime was reported in 2023 [1]. In [10], the behavior of two-photon states in non-Hermitian systems across the EP was studied. The EP was identified by analyzing variations in the Hong-Ou-Mandel interference with two indistinguishable input photons. The objective of the present work is to compute the EP in a PT-symmetric optical waveguide array with a quantum input state. To this end, two methods are employed: one based on the analysis of the system’s eigenvalues, and the other through the examination of the propagation patterns.
@@ 2 Theory
Figure 1: PT-symmetric system for (a) even number, and (b) odd, number of waveguides
@@
$[\hat{a}_{j},\hat{a}_{k}^{\dagger}]=\delta_{jk},\quad[\hat{a}_{j},\hat{a}_{k}]=0,\quad[\hat{a}_{j}^{\dagger},\hat{a}_{k}^{\dagger}]=0.$ (1)
If the propagation constant of the $j$-th waveguide is denoted by $\beta_{j}$ and the coupling coefficient between two adjacent waveguides is $c$ then the Hamiltonian describing the waveguide array is given by [13]:
@@
where the coupling between higher-order neighboring waveguides is neglected. The first term in equation (2) represents the transition of a photon from waveguide $j$ to $j+1$, while the second term corresponds to the reverse process. The final term represents the energy corresponding to each waveguide [13].
@@
$i\frac{\partial\hat{a}_{j}(z)}{\partial z}=c_{j-1}\hat{a}_{j-1}(z)+c_{j}\hat{a}_{j+1}(z)+\beta_{j}\hat{a}_{j}(z),$ (4)
@@
×$\quad\times G_{ik}(z)\,G_{jl}(z)\,\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle$ (10)
@@
$\begin{split}&\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=\frac{n}{2}\big{(}\delta_{np}\delta_{mp}+\delta_{nq}\delta_{mp}+\delta_{np}\delta_{mq}+\delta_{nq}\delta_{mq}\big{)},\\[6.0pt] &\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=\frac{n(n{-}1)}{2}\bigg{(}\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp}+\delta_{mq}\delta_{nq}\delta_{kq}\delta_{lq}\\ &\quad+\delta_{n,2}\big{(}\delta_{mp}\delta_{np}\delta_{kq}\delta_{lq}+\delta_{mq}\delta_{nq}\delta_{kp}\delta_{lp}\big{)}\bigg{)},\\[6.0pt] &I_{j}(z)=\frac{n}{2}\Big{(}|G_{jp}(z)|^{2}+|G_{jq}(z)|^{2}\\ &\quad+G^{*}_{jp}(z)\,G_{jq}(z)+G_{jp}(z)\,G^{*}_{jq}(z)\Big{)},\\[6.0pt] &g_{ji}^{(2)}(z)=\frac{n(n{-}1)}{2}\Big{(}|G_{jp}(z)G_{ip}(z)|^{2}+|G_{jq}(z)G_{iq}(z)|^{2}\\ &\quad+\delta_{n,2}\big{(}G^{*}_{jp}(z)G^{*}_{ip}(z)G_{jq}(z)G_{iq}(z)\\ &\qquad+G^{*}_{jq}(z)G^{*}_{iq}(z)G_{jp}(z)G_{ip}(z)\big{)}\Big{)}\end{split}$ (14)
@@
$I_{r}=\exp\left(-\alpha\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}\right)$ (15)
The determination is based on the spatial intensity distribution along the propagation length, with $\bar{I}_{1}$ and $\bar{I}_{2}$ representing the mean intensities over all waveguides in the first half ($0<z<\frac{z_{\text{max}}}{2}$) and second half ($\frac{z_{\text{max}}}{2}<z<z_{\text{max}}$) of the propagation length $z_{\text{max}}$, respectively. Besides, the parameters $\alpha$ and $\beta$ are two positive constants chosen to exhibit a significant change at the EP. In this study, the values $\alpha=\beta=10$ are employed. Before the EP, the energy exhibits oscillatory behavior, so $\bar{I}_{1}\simeq\bar{I}_{2}$. In this case, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ is approximately zero, which results in $I_{r}=1$. However, after the EP, as the propagation length increases, the energy grows exponentially, so $\bar{I}_{1}\ll\bar{I}_{2}$. Therefore, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ approaches one, and $I_{r}=0$. Subsequently, by plotting $I_{r}$ versus gain/loss factor $g$ and investigating the transition from one to zero in the $I_{r}$ criterion, the PT-symmetry breaking point $g_{c}$ can be identified.
A more accurate approach to locate the EP is to diagonalize the system Hamiltonian and calculate its eigenvalues. The EP is found at a critical gain/loss strength $g_{c}$ where the Hamiltonian’s eigenvalues shift from real to complex, which can be observed by plotting their imaginary parts versus $g$. The Hamiltonian (2) is defined in an infinite-dimensional Hilbert space. Consequently, determining its eigenvalues presents a significant analytical and computational challenge. Because the Hamiltonian (2) preserves the total photon number, as ensured by the commutation relation $[\hat{H},\hat{n}_{t}]=0$ with $\hat{n}_{t}=\sum_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j}$, the system’s dynamics can be analyzed within separate photon-number subspaces (zero, one, two photons, etc.), simplifying the problem.
@@
$\left(\sum_{j=1}^{N}\beta_{j}n_{j}^{\prime}\right)\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{N}^{\prime},n_{N}}$
$+c\sum_{j=1}^{N}\left\{\sqrt{n_{j}^{\prime}(n_{j+1}^{\prime}+1)}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}+1}\right.$
$\left.\delta_{n_{j+1}^{\prime},n_{j+1}-1}\dots\delta_{n_{N}^{\prime},n_{N}}\right.$
$+\sqrt{(n_{j}^{\prime}+1)n_{j+1}^{\prime}}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}-1}$
$\left.\delta_{n_{j+1}^{\prime},n_{j+1}+1}\dots\delta_{n_{N}^{\prime},n_{N}}\right\}$

### v2

# Intensity-Based Criterion for Determining Exceptional Point in Parity-Time (PT) Symmetric Coupled Array of Optical Waveguides

@@ Abstract
In this study, we investigated the propagation pattern and the site-to-site correlation function in a PT-symmetric waveguide array with different input quantum states. Recognizing the stark difference in propagation pattern before and after the PT symmetry-breaking point, we have developed a novel, straightforward intensity-based criterion to determine the exceptional point (EP). This new criterion shows excellent agreement with those obtained by directly computing the Hamiltonian’s eigenvalues. Given the computational complexity of Hamiltonian diagonalization, our proposed criterion provides a highly efficient and valuable alternative for identifying the PT symmetry-breaking point. Importantly, the proposed criterion is not restricted to the specific system studied here, but is generally applicable to a wide class of systems that can be described within the tight-binding framework.
@@ 1 Introduction
Over the past decade, discrete optical systems-particularly waveguide arrays-have played a crucial role in controlling the behavior of light [35]. Optical waveguide arrays consist of several single-mode waveguides arranged at specific distances. In this system, light is transferred between adjacent waveguides through evanescent coupling. The most effective method for manufacturing these waveguides is direct writing with femtosecond lasers [23]. Optical waveguide arrays represent one of the most advanced platforms for quantum information processing and quantum system simulations. This pivotal role is widely substantiated by extensive research, including experimental boson sampling [37], spatial entanglement in photonic lattices [6, 31], quantum walks of correlated photon pairs [29], generation of high-order single-photon W-states [14], experimental observation of N00N state Bloch oscillations [22], two-photon correlation and transport in disordered lattices [40] and anomalous photon transparency in non-Markovian coupled waveguides [32].
@@
The ability to control various parameters within waveguides offers a significant advantage for investigating PT-symmetric systems [11, 3, 20, 17, 16, 18, 12, 35]. PT-symmetry in quantum mechanics was first explored in 1998, following the discovery of Hamiltonians that remain invariant under both parity and time-reversal operations [30]. Notably, these non-Hermitian Hamiltonians can exhibit entirely real spectra [1]. PT-symmetry was introduced into optics due to the similarity between the paraxial wave equation and the Schrödinger equation, with the refractive index serving as the optical equivalent of the potential [8]. A PT-symmetric system in optics can be realized by adjusting the real and imaginary components of the refractive index [2]. For PT-symmetry to hold in optics, the real part of the refractive index needs to be spatially even, while its imaginary part must be odd function of position [27]. Since the imaginary part of the refractive index defines gain and loss, a PT-symmetric system can be achieved in optics by tuning these parameters accordingly. In a two-component optical system, real eigenvalues arise when the gain in one component compensates for the loss in the other. In contrast, in the weak coupling regime, if the gain from one component cannot fully compensate for the loss in the other, the system exhibits complex eigenvalues. This phase transition from real to complex spectra occurs at the EP [39, 19].
@@
A non-Hermitian Hamiltonian typically leads to complex eigenvalues, where the negative imaginary part represents a damping constant associated with the finite spectral linewidth of the resonance energy [38]. The real part of the eigenvalue determines this linewidth. In classical optical systems, the resonance frequencies and modes correspond to the eigenvalues and eigenstates of the Hamiltonian. Non-Hermiticity can arise due to absorption, gain, and radiation. When the Hamiltonian is non-Hermitian, i.e $\hat{H}\neq\hat{H}^{\dagger}$, its eigenstates lose their orthogonality with respect to the inner product used in the standard (Hermitian) formalism of quantum mechanics. However, this orthogonality can be restored by employing a new definition for the inner product, known as the c-product [26]. The non-orthogonality, under the standard inner product of quantum mechanics, is maximized at the EP, where the eigenstates of the Hamiltonian coalesce. In this scenario, the states become parallel to one another. These points are considered singularities within non-Hermitian systems. Near an EP, one or more eigenvalues and their corresponding eigenstates coalesce, leading to a phase transition. This transition, known as a PT-symmetry phase transition, results in the emergence of complex eigenvalues [39, 21].
@@
Another notable application in PT-symmetric optical systems is the achievement of zero group velocity at an exceptional point, where two optical modes coalesce. This effect, tunable across a wide range of frequencies and bandwidths, has been demonstrated in coupled waveguides with balanced gain and loss [10].
@@
Thompson et al. [36] examined intensity and site-to-site correlations in disordered optical waveguide arrays under Anderson localization, using non-classical (squeezed) and classical-like (coherent) light. However, their system did not exhibit PT symmetry. In 2018, the PT-symmetry breaking point of two curved, closely spaced PT-symmetric waveguides, characterized by a small period, was numerically computed using COMSOL Multiphysics software [42]. The investigation of boundary effects on propagation dynamics and site-to-site correlations within optical waveguide arrays in the quantum regime was reported in 2023 [31]. In [21], the behavior of two-photon states in non-Hermitian systems across the EP was studied. The EP was identified by analyzing variations in the Hong-Ou-Mandel interference with two indistinguishable input photons. The objective of the present work is to compute the EP in a PT-symmetric optical waveguide array with a quantum input state. To this end, two methods are employed: one based on the analysis of the system’s eigenvalues, and the other through the examination of the propagation patterns. Furthermore, we show that the proposed intensity-based criterion is in excellent quantitative agreement with the conventional eigenvalue-based determination of the PT-symmetry breaking point. Considering that the direct diagonalization of the Hamiltonian, especially in quantum regimes characterized by exponentially growing Hilbert spaces, is computationally intensive, the present approach offers a highly efficient and scalable alternative. Notably, the scope of this framework extends well beyond the specific photonic system studied here, encompassing a wide range of Schrödinger-type dynamics that admit a tight-binding representation, such as paraxial wave propagation in optical and mechanical settings, quantum evolution in spin chains, population transfer in multilevel configurations, and condensed-matter systems.
@@ 2 Theory
$\hat{a}_{j},\hat{a}_{k}^{\dagger}]=\delta_{jk},\quad[\hat{a}_{j},\hat{a}_{k}]=0,\quad[\hat{a}_{j}^{\dagger},\hat{a}_{k}^{\dagger}]=0.$ (1)
Figure 1: PT-symmetric system for (a) even number, and (b) odd, number of waveguides. The waveguides exhibit cylindrical symmetry and are located at equal distances. The colors red, green, and gray correspond to gain (G), loss (L), and neutral (N) conditions, respectively. In these figures, the yellow arrow shows how light enters the system.
If the propagation constant of the $j$-th waveguide is denoted by $\beta_{j}$ and the coupling coefficient between two adjacent waveguides is $c$ then, in the tight-binding approximation, the Hamiltonian describing the waveguide array is given by [25]:
@@
where $c$ is the coupling coefficient between nearest-neighbor waveguides, assumed to be identical for all pairs. Moreover, the coupling between higher-order neighboring waveguides is neglected. The first term in equation (2) represents the transition of a photon from waveguide $j$ to $j+1$, while the second term corresponds to the reverse process. The final term represents the energy corresponding to each waveguide [25].
It is important to note that the geometrical configuration of waveguide arrays used in reported experiments explicitly confirms the necessary conditions for applying the tight-binding approximation and evanescent nearest-neighbor coupling. The distance between the centers of adjacent waveguides is typically significantly larger than the excitation wavelength. Therefore, the coupling between adjacent waveguides occurs solely through the overlap of evanescent fields. Consequently, the physics of these evanescently coupled waveguide structures can be accurately described by the tight-binding approximation. This applicability is further validated by the high degree of similarity demonstrated in numerous studies between theoretical predictions (based on the tight-binding model) and experimental results [37, 6, 29, 14, 22, 40, 32, 11, 3, 20, 35].
@@
It should be noted that the ideal treatment of gain and loss effects in a full quantum optical framework involves the Heisenberg-Langevin equations [13] (in the Heisenberg picture) or solving the Master equation [31, 5] (in the Schrödinger picture). However, for simplicity, our study, following the methodology of references [40, 15, 41, 43, 9, 28, 4], adopts the standard non-Hermitian Hamiltonian formalism, where gain and loss are incorporated as imaginary potential terms. The non-Hermitian Hamiltonian captures the average, or mean-field, effect of the environment on the subsystem’s dynamics. While it does not account for quantum noise and fluctuations in the same way the full Heisenberg-Langevin or Master equation does, it has been shown to provide an excellent approximation for studying phenomena that are primarily determined by the spectrum of the Hamiltonian, such as exceptional points and their associated dynamics. This approach, which is extensively used and validated in numerous publications on PT-symmetric optics and waveguide arrays (Refs. [40, 15, 41, 43, 9, 28, 4]), providing a balance between theoretical rigor and computational feasibility.
@@
$i\frac{\partial\hat{a}_{j}(z)}{\partial z}=c\,\hat{a}_{j-1}(z)+c\,\hat{a}_{j+1}(z)+\beta_{j}\hat{a}_{j}(z),$ (4)
@@
Intensity, provides information about the probability of detecting photon at a specific location. This metric is useful for understanding how optical energy is distributed within the array. However, intensity alone cannot reveal deeper quantum aspects of the system, such as spatial entanglement or photon statistics. In essence, intensity shows the average collective behavior of photons.
@@
$\quad\times G_{ik}(z)\,G_{jl}(z)\,\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle.$ (10)
The second-order correlation function can reveal whether photons behave randomly, exhibit photon bunching, or photon antibunching. This information is essential for understanding the quantum nature of light and investigating spatial entanglement between photons, which is not observable solely from intensity [6, 29, 40]. Photons can travel between distant waveguides through a series of sequential nearest-neighbor hops, thereby causing the optical fields at non-adjacent locations to become correlated. Therefore, while photons are not directly coupled between distant waveguides, they can interact indirectly, via intermediate waveguides, leading to correlations in their detection. The site-to-site correlation function precisely measures this spatial correlation.
@@
$\begin{split}&\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=\frac{n}{2}\big(\delta_{np}\delta_{mp}+\delta_{nq}\delta_{mp}+\delta_{np}\delta_{mq}+\delta_{nq}\delta_{mq}\big),\\[6.0pt] &\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=\frac{n(n-1)}{2}\bigg(\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp}+\delta_{mq}\delta_{nq}\delta_{kq}\delta_{lq}\\ &\qquad+\delta_{n,2}\big(\delta_{mp}\delta_{np}\delta_{kq}\delta_{lq}+\delta_{mq}\delta_{nq}\delta_{kp}\delta_{lp}\big)\bigg),\\[6.0pt] &I_{j}(z)=\frac{n}{2}\Big(|G_{jp}(z)|^{2}+|G_{jq}(z)|^{2}\\ &\qquad+G^{*}_{jp}(z)\,G_{jq}(z)+G_{jp}(z)\,G^{*}_{jq}(z)\Big),\\[6.0pt] &g_{ji}^{(2)}(z)=\frac{n(n{-}1)}{2}\Big(|G_{jp}(z)G_{ip}(z)|^{2}+|G_{jq}(z)G_{iq}(z)|^{2}\\ &\qquad+\delta_{n,2}\big(G^{*}_{jp}(z)G^{*}_{ip}(z)G_{jq}(z)G_{iq}(z)\\ &\hskip 18.49988pt+G^{*}_{jq}(z)G^{*}_{iq}(z)G_{jp}(z)G_{ip}(z)\big)\Big)\end{split}$ (14)
@@
$I_{r}=\exp\left(-\gamma\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}\right)$ (15)
The determination is based on the spatial intensity distribution along the propagation length, with $\bar{I}_{1}$ and $\bar{I}_{2}$ representing the mean intensities over all waveguides in the first half ($0<z<\frac{z_{\text{max}}}{2}$) and second half ($\frac{z_{\text{max}}}{2}<z<z_{\text{max}}$) of the propagation length $z_{\text{max}}$, respectively. Besides, the parameters $\alpha$ and $\beta$ are two positive constants chosen to exhibit a significant change at the EP. In practice, the choice $\gamma=\beta=10$ provides a robust and reliable characterization of the transition, and the criterion is not highly sensitive to the precise values of these parameters. To further clarify the validity and underlying mechanism of the proposed criterion, we examine its behavior in the two distinct regimes of unbroken and broken PT symmetry. Before the EP, the energy exhibits oscillatory behavior, so $\bar{I}_{1}\simeq\bar{I}_{2}$. In this case, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ is approximately zero, which results in $I_{r}=1$. However, after the EP, as the propagation length increases, the energy grows exponentially, so $\bar{I}_{1}\ll\bar{I}_{2}$. Therefore, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ approaches one, and $I_{r}=0$. These distinct behaviors lead to clearly different values of the proposed measure in the two regimes, thereby enabling it to serve as a reliable indicator for identifying the transition point. Subsequently, by plotting $I_{r}$ versus gain/loss factor $g$ and investigating the transition from one to zero in the $I_{r}$ criterion, the PT-symmetry breaking point $g_{c}$ can be identified.
A more accurate approach to locate the EP is to diagonalize the system Hamiltonian and calculate its eigenvalues. The EP is found at a critical gain/loss strength $g_{c}$ where the Hamiltonian’s eigenvalues shift from real to complex, which can be observed by plotting their imaginary parts versus $g$. Because of the bosonic statistics of photons, there is no inherent limit to the number of photons residing in each waveguide. Therefore, the Hamiltonian (2) is defined in an infinite-dimensional Hilbert space. Consequently, determining its eigenvalues presents a significant analytical and computational challenge. Moreover, if the maximum number of photons per waveguide is restricted to $N_{p}$, the dimension of the aforementioned Hilbert space becomes $(N_{p}+1)^{N}$, which exponentially increases with number of waveguides, N. The exponential growth of the Hilbert space dimension with the number of waveguides, and the associated computational complexity of its diagonalization, highlight the importance of the criterion introduced in Eq. (15) for determining the EP. This is because calculating this criterion does not require Hamiltonian diagonalization; instead, the EP can be determined by solving the $N^{2}$ coupled linear differential equations in Eq. (6) and computing the spatial evolution of the intensity. In this research, because the Hamiltonian (2) preserves the total photon number, as ensured by the commutation relation $\hat{H},\hat{n}_{t}]=0$ with $\hat{n}_{t}=\sum_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j}$, the system’s dynamics can be analyzed within separate photon-number subspaces (zero, one, two photons, etc.), simplifying the problem.
@@
$\hskip-184.9429pt\left(\sum_{j=1}^{N}\beta_{j}n_{j}^{\prime}\right)\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{N}^{\prime},n_{N}}$
$\hskip-184.9429pt+c\sum_{j=1}^{N}\left\{\sqrt{n_{j}^{\prime}(n_{j+1}^{\prime}+1)}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}+1}\right$.
$\hskip-184.9429pt\left.\delta_{n_{j+1}^{\prime},n_{j+1}-1}\dots\delta_{n_{N}^{\prime},n_{N}}\right$.
$\hskip-184.9429pt+\sqrt{(n_{j}^{\prime}+1)n_{j+1}^{\prime}}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}-1}$
$\hskip-184.9429pt\left.\delta_{n_{j+1}^{\prime},n_{j+1}+1}\dots\delta_{n_{N}^{\prime},n_{N}}\right\}$